# Alleviating Distribution Shift in Synthetic Data for Machine Translation Quality Estimation

### v2

@@ Appendix D Time Analysis
Higher-quality pseudo-data can accelerate training convergence. To evaluate the convergence efficiency of different methods, we plot the learning curve on the WMT2022 validation set in Figure 4. For both sentence-level and word-level tasks, CBSQE demonstrates faster convergence compared to MQMQE, ultimately achieving the highest performance.
@@ Appendix F Amount of synthetic Data
We examine the impact of data size on the performance of ADSQE. To this end, we conduct a series of experiments utilizing varying amounts of parallel sentence pairs from the WMT23 EN-DE dataset. As illustrated in Figure 5, for all the tasks, pseudo-data methods demonstrate significant performance improvements as the data size increases. Notably, this upward trend persists and does not exhibit convergence until reaching a scale of 5M parallel pairs.

### v3

@@ Appendix D Time Analysis
Higher-quality synthetic data can accelerate training convergence. To evaluate the convergence efficiency of different methods, we plot the learning curve on the WMT2022 validation set in Figure 4. For both sentence-level and word-level tasks, CBSQE demonstrates faster convergence compared to MQMQE, ultimately achieving the highest performance.
@@ Appendix F Amount of Synthetic Data
We examine the impact of data size on the performance of DCSQE. To this end, we conduct a series of experiments utilizing varying amounts of parallel sentence pairs from the WMT23 EN-DE dataset. As illustrated in Figure 5, for all the tasks, synthetic data methods demonstrate significant performance improvements as the data size increases. Notably, this upward trend persists and does not exhibit convergence until reaching a scale of 5M parallel pairs.
@@
Appendix G Threshold Sensitivity
In DCSQE, the thresholds can be easily acquired by small size validation sets for each language pair which are available in most real world scenarios. For a given language pair, these thresholds exhibit robustness across different datasets.
In our experiments, we apply the same thresholds derived from the WMT2022 validation dataset to both the WMT2022 and WMT2023 test sets. As shown in Table 2, DCSQE still achieves consistent performance across different test sets for both EN-DE and ZH-EN language pairs. Notably, there exist significant differences between WMT2022 and WMT2023 test sets as discussed in Blain et al. (2023).
Appendix H Completeness of Error Modeling
Alignment between translations and references is generally effective in identifying most translation errors, as these typically appear as discrepancies with the references. Our manual inspection of 100 randomly sampled synthetic instances, does not reveal any undetected errors (i.e., false-positives labels). This observation is consistent with the finding in CBSQE Geng et al. (2023b), which reports that the TER tool frequently produces false-negative labels but rarely false positives.
To more comprehensively validate the comprehensiveness of error modeling in DCSQE, we compare the error detection accuracy of models without/with DCSQE synthetic data pre-training across various error categories. Since error category labels are only available in training data, we use the WMT2024 training set [^10] https://github.com/google/wmt-mqm-human-evaluation/tree/main/generalMT2023 which is never exposed during our model training. The experimental results demonstrate significant improvements in nearly all error categories (including semantic errors), as shown in the Table 13. This finding confirms that our synthetic data encompasses diverse error types.